On a research it was found that 47% of residents travelers were women and 53% men. A random sample of 500 travelers on a large airline revealed that of those 500, 263 were women. Using a=0.05 compute CV, and TV, and state the hypotheses and your conclusion.


1
Expert's answer
2021-05-07T09:10:04-0400

Null hypothesis "H_0:p=0.47."

Alternative hypothesis "H_a:p\ne 0.47."

CV=1.96.

TV"=\frac{\frac{263}{500}-0.47}{\sqrt{\frac{0.47(1-0.47)}{500}}}=2.51."

Since TV is greater than CV, reject the null hypothesis.

There is a sufficient evidence that percent of women travelers is different from 47%.